# Koopman for stochastic dynamics: error bounds for kernel extended dynamic mode decomposition

### v1

@@ 2.2. Recap on deterministic systems: kernel EDMD and error bound
    for a suitable rate $r>0$ and $C\geq 0$ using existing results from kernel-based approximation theory, see e.g. [Wend04, Section 11.2].
The invariance property (1) is a subtle detail that has to be verified rigorously and cannot be taken for granted. In particular, there are RKHS for which $\mathcal{K}\in L(\mathbb{H},\mathbb{H})$ implies that the underlying dynamics is affine linear, such as the Gaussian RKHS, see [GonzAbud23]. For (fractional) Sobolev spaces $H^{s}$ (see Section 3.2 for an introduction), however, which are RKHS if $s>n/2$ and are generated by Wendland or Matérn kernels (see below for a precise definition), Koopman invariance was proven in [KohnPhil24, Theorem 4.2] in the deterministic setting. The central assumption is that the flow is smooth enough, such that the Koopman composition operator preserves Sobolev regularity. On the other hand, invariance of some RKHS under the Koopman operator is sufficient for boundedness, as shown in [KohnPhil24, Lemma 3.1] for the deterministic case, and analogously in Proposition 2.4 below for the stochastic case.
@@
Since $\mathbb{H}\subset C_{b}(\Omega)$, the corresponding kernel is bounded. Let $f_{n},f,g\in\mathbb{H}$ such that $f_{n}\to f$ and $\mathcal{K}f_{n}\to g$ in $\mathbb{H}$. Recall the canonical features $\Phi_{x}$ for $x\in\Omega$. Then, the reproducing property (8) yields
$g(x)$ $=\langle g,\Phi_{x}\rangle_{\mathbb{H}}=\lim_{n\to\infty}\langle\mathcal{K}f_{n},\Phi_{x}\rangle_{\mathbb{H}}=\lim_{n\to\infty}\mathcal{K}f_{n}(x)=\lim_{n\to\infty}\int f_{n}(y)\,d\rho_{x}(y).$
@@
To sum up, the assumption of Theorem 2.3 is met for Sobolev RKHS, and error bounds of the projection error (14) are readily available from the literature for the generating kernels, leading to a pointwise and deterministic error bound for EDMD for deterministic systems.
@@ 2.3. Kernel EDMD for stochastic systems
In contrast to the deterministic case, the stochastic Koopman operator is defined by means of conditional expectations. These are usually not readily available in applications with unknown dynamics such that the analogue to the regression problem (6) is not directly solvable if only observations are available. More precisely, to implement the counterpart of the approximate identity (5), i.e.,
$K\Psi_{\mathcal{X}}\approx\Psi_{\mathcal{Y}}=(\mathbb{E}[\Psi_{i}(Y)\mid X=x_{j}])_{(i,j)\in[M]\times[d]},$ (15)
or the corresponding regression problem (6), the data matrix $\Psi_{\mathcal{Y}}$ has to be further approximated. To this end, we build upon the rich theory of Monte Carlo methods to perform an entry-wise approximation, where we draw $m_{j}\in\mathbb{N}$ i.i.d. samples $y_{j}^{(l)}\sim\rho_{x_{j}}$, $l=1,\dots,m_{j}$, and invoke the law of large numbers such that
@@
For the kernel case, we then have, in view of $\Psi_{i}=k(x_{i},\cdot)$,
$K\mathbf{K}_{\mathcal{X}}\approx\mathbf{K}_{\mathcal{X},\mathcal{Y}}^{\mathrm{MC}}\qquad\mathrm{with}\qquad(\mathbf{K}_{\mathcal{X},\mathcal{Y}}^{\mathrm{MC}})_{ij}=\frac{1}{m_{j}}\sum_{l=1}^{m_{j}}k(x_{i},y_{j}^{(l)})$ (17)
@@
$\widehat{\mathcal{K}}^{\mathrm{MC}}f=f_{\mathcal{X}}^{\top}\,\mathbf{K}_{\mathcal{X}}^{-1}\,\mathbf{K}_{\mathcal{X},\mathcal{Y}}^{\mathrm{MC}}\,\mathbf{K}_{\mathcal{X}}^{-1}\,\mathbf{k}_{\mathcal{X}}$ (18)
which mirrors the deterministic approximant from (13). However, the entry-wise Monte Carlo approximation $\mathbf{K}_{\mathcal{X},\mathcal{Y}}^{\mathrm{MC}}$ introduces a further estimation error since (16) only holds approximately. In the ideal case where $\Psi_{\mathcal{Y}}=\mathbf{K}_{\mathcal{X},\mathcal{Y}}$ is known exactly, this estimation error vanishes, and the error bound from Theorem 2.3 holds.
@@
Another possible design choice more in line with the covariance operator approach circumvents the estimation of estimating the expected value under certain assumptions on the underlying stochastic system and sampling scheme (see [NuskPeit23, PhilSchaWort2025] for details). In the setting of a governing stochastic differential equation, the authors either sample the design points in $\mathcal{X}$ from the invariant measure $\mu$ or assume ergodicity and sample from a single long trajectory $x_{0},x_{1},\dots,x_{d}$. In this paper, we pursue the first approach. Future work utilizing the second approach could be the subject of further work and will be discussed later.
@@ 2.4. Main contribution: error analysis of kEDMD for stochastic systems
Let $\mathbb{H}$ be an RKHS of functions over $\Omega$, and assume $\mathcal{K}\in L(\mathbb{H},\mathbb{H})$ and $\mathcal{K}\in L(C_{b}(\Omega),C_{b}(\Omega))$. Furthermore, for $\mathcal{X}=\{x_{1},\ldots,x_{d}\}$, let $m_{j}=m\in\mathbb{N}$ i.i.d. samples from every $\rho_{x_{j}}$, $j\in\{1,\dots,d\}$ be given. Then, for every $\varepsilon>0$ and for all $f\in\mathbb{H}$, $\|f\|_{\mathbb{H}}\leq 1$, we have the pointwise error bound
$\|(\mathcal{K}-\widehat{\mathcal{K}}^{\mathrm{MC}})f\|_{\infty}\leq\big[1+\|\mathcal{K}\|_{\mathbb{H}\to\mathbb{H}}\big]\|\mathcal{I}-S_{\mathcal{X}}\|_{\mathbb{H}\to C_{b}}+\varepsilon\cdot\|k\|_{\infty}^{1/2}\sqrt{\max_{v\in\mathbbm{1}}v^{\top}\mathbf{K}_{\mathcal{X}}^{-1}v}$ (19)
with probability at least $1-2d\exp\left(-\frac{m\varepsilon^{2}}{2\|k\|_{\infty}}\right)$.
The proof of the above theorem consists of two parts. First, the finite number of data points and observables given by $\mathcal{X}\subset\Omega$ introduces a deterministic error. Second, we can only estimate the entries of the “propagation matrix” $\mathbf{K}_{\mathcal{X},\mathcal{Y}}^{\mathrm{MC}}$ which are conditional expected values. Our estimator $\mathbf{K}_{\mathcal{X},\mathcal{Y}}^{\mathrm{MC}}$ is random itself, resulting in a probabilistic estimation error. Thus, we split the approximation error as follows:
@@
The probabilistic error is controlled by the accuracy of the entry-wise Monte Carlo estimator (5), i.e., the number $m$ of i.i.d. samples from $\rho_{x_{j}}$. Besides, the two main components in Theorem 2.6 are the operator norms (which are assumed to be finite) and the projection error governed by $\mathcal{I}-S_{\mathcal{X}}$. These two components strongly rely on the chosen kernel and thus deserve separate treatments which constitute the second and third main result of this work.
Boundedness of the Koopman operator. Invariance of reproducing kernel Hilbert spaces for Koopman operators is a non-trivial issue. Hence, in Section 3 we provide characterizations and sufficient conditions for the invariance of Sobolev spaces. Here, we present the most important and practically relevant result. Its proof follows from the more general result of Theorem 3.2.
@@
If the transition kernel $\rho_{x}$ is given by a density $p:\Omega\times\Omega\to[0,\infty)$, i.e.,
$\rho_{x}(\Delta)=\int_{\Delta}p(x,y)\,dy\qquad\forall x\in\Omega,\,\Delta\in\mathcal{B}_{\Omega},$
that satisfies
@@
for some constant $C>0$ and $\ell\in\mathbb{N}$, then $\mathcal{K}\in L(H^{s},H^{s})$ for all $s\in(0,\ell]$.
We point out that Koopman invariance, as shown in [KohnPhil24] for suitably chosen reproducing kernel Hilbert spaces, was extended to stochastic systems in the preprint [OlguOsseRami2025], under the stronger regularity assumption $p\in C^{m}(\Omega\times\Omega)$ for the transition density. On the one hand, we generalize this result by significantly weakening the assumptions. On the other hand, we also provide necessary conditions and, thus, a complete characterization – even without assuming that a density for the transition kernel $\rho_{x}$ exists.
Projection error bounds in terms of the fill distance. In this work, we apply our results to two popular choices of kernels, Wendland kernels [Wend04] and Matérn kernels [BerlThom2004] which are in particular consistent with the above mentioned invariance as they generate Sobolev spaces (see [Adams2003] and Section 3.2 for an introduction). We briefly present the error bounds for Wendland kernels, as the bounds for Matérn kernels (see Appendix C for an introduction) are structurally very similar.
The Wendland kernel $k_{n,l}:\mathbb{R}^{n}\times\mathbb{R}^{n}\to\mathbb{R}$ is bounded, continuous and strictly positive-definite (see [Wend04] and Appendix B for details). For $\sigma_{n,l}=\frac{n+1}{2}+l$, it generates the (fractional) Sobolev space $H^{\sigma_{n,l}}(\mathbb{R}^{n})$ as RKHS with equivalence of the standard Sobolev norm and the induced (native) RKHS norm. Restricted to a bounded domain $\Omega\subset\mathbb{R}^{n}$ with Lipschitz boundary, as is of interest in the setting of this paper, the Wendland kernel generates the corresponding Sobolev space $H^{\sigma_{n,l}}(\Omega)$ [KohnPhil24, Theorem 4.1]. Results from kernel approximation theory allow for a more explicit bound on the projection error in terms of the fill distance.
@@
Let $l\in\mathbb{N}$, define $\sigma_{n,l}=\frac{n+1}{2}+l$ and denote by $\mathbb{H}$ the Sobolev space $H^{\sigma_{n,l}}(\Omega)$ induced by the Wendland kernel $k_{n,l}$ as RKHS. Assume that $\mathcal{K}\in L(\mathbb{H},\mathbb{H})$ and $\mathcal{K}\in L(C_{b}(\Omega),C_{b}(\Omega))$. Then, for every $\varepsilon>0$ and all $f\in\mathbb{H}$ with $\|f\|_{\mathbb{H}}\leq 1$, we have the pointwise error bound
$\|(\mathcal{K}-\mathcal{K}^{\mathrm{MC}})f\|_{\infty}\leq$ $C_{1}\,(1+\|\mathcal{K}\|_{\mathbb{H}\to\mathbb{H}})\cdot h_{\mathcal{X}}^{l+1/2}+\varepsilon\cdot C_{2}\cdot\sqrt{\max_{v\in\mathbbm{1}}v^{\top}\mathbf{K}_{\mathcal{X}}^{-1}v}$ (21)
with probability at least $1-2d\exp\left(-\frac{m\varepsilon^{2}}{2C_{2}^{2}}\right)$, with constants $C_{1},C_{2}>0$ only depending on the kernel $k_{n,l}$.
Note that the bounds of Theorems 2.6 and 2.8 depend on the smallest eigenvalue of the kernel matrix $\mathbf{K}_{\mathcal{X}}$. Tikhonov regularization is a means to control the latter, for the price of introducing a deterministic bias Section 4 addresses the impact of regularization on the deterministic (Proposition 4.5), probabilistic (Corollary 4.9) and total error bounds (Corollary 4.12).

### v2

@@ 2.2. Recap on deterministic systems: kernel EDMD and error bound
    for a suitable rate $r>0$ and $C\geq 0$ using existing results from kernel-based approximation theory, see e.g. [50, Section 11.2].
The invariance property (1) is a subtle detail that has to be verified rigorously and cannot be taken for granted. In particular, there are RKHS for which $\mathcal{K}\in L(\mathbb{H},\mathbb{H})$ implies that the underlying dynamics is affine linear, such as the Gaussian RKHS, see [15]. For (fractional) Sobolev spaces $H^{s}$ (see Section 3.2 for an introduction), however, which are RKHS if $s>n/2$ and are generated by Wendland or Matérn kernels (see below for a precise definition), Koopman invariance was proven in [24, Theorem 4.2] in the deterministic setting. The central assumption is that the flow is smooth enough, such that the Koopman composition operator preserves Sobolev regularity. On the other hand, invariance of a RKHS under the Koopman operator is sufficient for boundedness, as shown in [24, Lemma 3.1] for the deterministic case, and analogously in Proposition 2.4 below for the stochastic case.
@@
Since $\mathbb{H}\subset C_{b}(\Omega)$, the corresponding kernel is bounded. Let $f_{n},f,g\in\mathbb{H}$ such that $f_{n}\to f$ and $\mathcal{K}f_{n}\to g$ in $\mathbb{H}$. Recall the canonical features $\Phi_{x}$ at $x\in\Omega$. Then, the reproducing property (9) yields
$g(x)$ $=\langle g,\Phi_{x}\rangle_{\mathbb{H}}=\lim_{n\to\infty}\langle\mathcal{K}f_{n},\Phi_{x}\rangle_{\mathbb{H}}=\lim_{n\to\infty}\mathcal{K}f_{n}(x)=\lim_{n\to\infty}\int_{\Omega}f_{n}(y)\,d\rho_{x}(y).$
@@
To sum up, the assumption of Theorem 2.3 is met for Sobolev RKHS, and bounds on the projection error in the form (15) are well-known for many generating kernels, leading to a pointwise and deterministic error bound for deterministic systems.
@@ 2.3. Kernel EDMD for stochastic systems
In constrast, the stochastic Koopman operator propagates conditional expectations of observables. Here, the counterpart of the approximate identity (6) takes the form
$K\Psi_{\mathcal{X}}\approx\Psi_{\mathcal{Y}}=\big(\mathbb{E}[\Psi_{i}(Y)\mid X=x_{j}]\big)_{i,j=1}^{M,d}.$ (16)
Consequently, for the finite-dimensional approximation $K\in\mathbb{R}^{M\times M}$ in (16), setting
$Z^{+}:=\Psi(Y)\approx KZ,\qquad Z:=\Psi(X),$
defines a linear surrogate system in feature space. The surrogate system describes the evolution of the conditional mean of the lifted observables, i.e., $\mathbb{E}[Z^{+}\,|\,X]\approx KZ=K\Psi(X)$. Here, probability enters through the conditional expectation: the surrogate system approximates the expected evolution of the lifted observables, while stochastic fluctuations around this mean remain. However, the conditional expectations in (16) are usually not readily available in applications with unknown dynamics, such that the analogue to the regression problem (7) cannot be directly solved. Instead, the data matrix $\Psi_{\mathcal{Y}}$ has to be approximated from observations of the stochastic system. For this, we build upon the rich theory of Monte Carlo methods to perform an entry-wise approximation. We draw $m_{j}\in\mathbb{N}$ i.i.d. samples $y_{j}^{(l)}\sim\rho_{x_{j}}$, $l=1,\dots,m_{j}$, and invoke the law of large numbers such that
@@
For the kernel case, we then have, in view of $\Psi_{i}=\text{\sf{k}}(x_{i},\cdot)$,
$K\mathbf{K}_{\mathcal{X}}\approx\mathbf{K}_{\mathcal{X},\mathcal{Y}}^{\mathrm{MC}}\qquad\mathrm{with}\qquad(\mathbf{K}_{\mathcal{X},\mathcal{Y}}^{\mathrm{MC}})_{ij}=\frac{1}{m_{j}}\sum_{l=1}^{m_{j}}\text{\sf{k}}(x_{i},y_{j}^{(l)})$ (18)
@@
$\widehat{\mathcal{K}}^{\mathrm{MC}}f=f_{\mathcal{X}}^{\top}\,\mathbf{K}_{\mathcal{X}}^{-1}\,\mathbf{K}_{\mathcal{X},\mathcal{Y}}^{\mathrm{MC}}\,\mathbf{K}_{\mathcal{X}}^{-1}\,\mathbf{k}_{\mathcal{X}},$ (19)
mirroring the deterministic approximant from (14). However, the entry-wise Monte Carlo approximation $\mathbf{K}_{\mathcal{X},\mathcal{Y}}^{\mathrm{MC}}$ introduces a further estimation error since (17) only holds approximately. In the ideal case where $\Psi_{\mathcal{Y}}=\mathbf{K}_{\mathcal{X},\mathcal{Y}}$ is known exactly, this estimation error vanishes, and the error bound from Theorem 2.3 holds.
@@
Another possible design choice more in line with the covariance operator approach circumvents the estimation of the expected value in (17) under certain assumptions on the underlying stochastic system and sampling scheme (see [36, 41] for details). In the setting of a governing stochastic differential equation, the authors either sample the design points in $\mathcal{X}$ from the invariant measure $\mu$ or assume ergodicity and sample from a single long trajectory $x_{0},x_{1},\dots,x_{d}$. In this paper, we pursue the Monte Carlo approach illustrated above, which requires weaker assumptions on the system but relies on i.i.d. samples.
@@ 2.4. Main contribution: error analysis of kEDMD for stochastic systems
Let $\mathbb{H}$ be an RKHS of functions over $\Omega$, and assume $\mathcal{K}\in L(\mathbb{H},\mathbb{H})$ and $\mathcal{K}\in L(C_{b}(\Omega),C_{b}(\Omega))$. Furthermore, for $\mathcal{X}=\{x_{1},\ldots,x_{d}\}\subset\Omega$, let $m_{j}=m\in\mathbb{N}$ i.i.d. samples from every $\rho_{x_{j}}$, $j\in\{1,\dots,d\}$, be given. Then, for every $\varepsilon>0$ satisfying $m\geq\frac{2\|\text{\sf{k}}\|_{\infty}}{\varepsilon^{2}}\log(2d)$ and for all $f\in\mathbb{H}$, $\|f\|_{\mathbb{H}}\leq 1$, we have the pointwise error bound
$\big\|\big(\mathcal{K}-\widehat{\mathcal{K}}^{\mathrm{MC}}\big)f\big\|_{\infty}\leq\big[1+\|\mathcal{K}\|_{\mathbb{H}\to\mathbb{H}}\big]\|\mathcal{I}-S_{\mathcal{X}}\|_{\mathbb{H}\to C_{b}}+\varepsilon\cdot\|\text{\sf{k}}\|_{\infty}^{1/2}\sqrt{\max_{v\in\mathbbm{1}}v^{\top}\mathbf{K}_{\mathcal{X}}^{-1}v}$ (20)
with probability at least $1-2d\exp\left(-\frac{m\varepsilon^{2}}{2\|\text{\sf{k}}\|_{\infty}}\right)$, where $\mathbbm{1}=\{v\in\mathbb{R}^{d}:|v_{i}|=1\}$.
The proof of the above theorem (cf. Section 4.3) consists of two parts. First, the finite number of data points and observables given by $\mathcal{X}\subset\Omega$ introduces a deterministic error. Second, we can only estimate the entries of the “propagation matrix” $\mathbf{K}_{\mathcal{X},\mathcal{Y}}^{\mathrm{MC}}$ which are conditional expected values. Our estimator $\mathbf{K}_{\mathcal{X},\mathcal{Y}}^{\mathrm{MC}}$ is random itself, resulting in a probabilistic estimation error. Thus, we split the approximation error as follows:
@@
In Theorem 2.6, the parameter $m$ is chosen sufficiently large (depending on $\varepsilon$ and $d$) to ensure that the stated lower bound on the probability is nonnegative. More precisely, for fixed $d$ and $\varepsilon>0$, choosing $m\geq\frac{2\|\text{\sf{k}}\|_{\infty}}{\varepsilon^{2}}\log(2d)$ guarantees that the probability bound is meaningful. In practice, the parameters $d$ and $m$ play different roles: the number of design points $d$ controls the deterministic approximation error, while $m$ determines the accuracy of the Monte Carlo estimation. For a prescribed accuracy level $\varepsilon$, one therefore chooses $d$ large enough to control the deterministic error and then selects $m=m(d,\varepsilon)$ sufficiently large to achieve the desired error bound with high probability. We refer the reader to Section 5.1 and Appendix D for further discussion on the interplay between $d$, $\varepsilon$, and $m$.
The assumption $\mathcal{K}\in L(C_{b}(\Omega),C_{b}(\Omega))$ is a regularity condition on the transition kernel $\rho_{x}$, which is characterized and discussed in Section 4.1, and is independent of the continuity of the RKHS kernel k defining $\mathbb{H}$. Besides, the two main components in Theorem 2.6 are the operator norms (which are assumed to be finite) and the projection error governed by $\mathcal{I}-S_{\mathcal{X}}$. These two components strongly rely on the chosen kernel and thus deserve separate treatments which constitute the second and third main result of this work.
Boundedness of the Koopman operator. Invariance of reproducing kernel Hilbert spaces under Koopman operators is a non-trivial issue. Hence, in Section 3 we provide characterizations and sufficient conditions for the invariance of Sobolev spaces. Here, we present the most important and practically relevant result, which follows from the more general result of Theorem 3.2.
@@
If the transition kernel $\rho_{x}$ is given by a density $p:\Omega\times\Omega\to[0,\infty)$, i.e., $\rho_{x}(\Delta)=\int_{\Delta}p(x,y)\,dy$ for all $x\in\Omega$ and $\Delta\in\mathcal{B}_{\Omega}$, that satisfies
@@
for some constant $C>0$ and $\ell\in\mathbb{N}$, then $\mathcal{K}\in L(H^{s}(\Omega),H^{s}(\Omega))$ for all $s\in(0,\ell]$.
We point out that Koopman invariance, as shown in [24] for suitably chosen reproducing kernel Hilbert spaces, was extended to stochastic systems in the preprint [38], under the stronger regularity assumption $p\in C^{m}(\Omega\times\Omega)$ for the transition density. On the one hand, we generalize this result by significantly weakening the assumptions. On the other hand, we also provide necessary conditions and, thus, a complete characterization – even without assuming that a density for the transition kernel $\rho_{x}$ exists.
Projection error bounds in terms of the fill distance. In this work, we apply our results to two popular choices of kernels, Wendland kernels [50] and Matérn kernels [6] which are in particular consistent with the above mentioned invariance as they generate Sobolev spaces (see [1] and Section 3.2 for an introduction). We briefly present the error bounds for Wendland kernels, as the bounds for Matérn kernels (see Appendix C for an introduction) are structurally very similar.
The Wendland kernel $\text{\sf{k}}_{n,l}:\mathbb{R}^{n}\times\mathbb{R}^{n}\to\mathbb{R}$ is bounded, continuous and strictly positive-definite (see [50] and Appendix B for details). For $\sigma_{n,l}=\frac{n+1}{2}+l$, it generates the (fractional) Sobolev space $H^{\sigma_{n,l}}(\mathbb{R}^{n})$ as RKHS with equivalence of the standard Sobolev norm and the induced (native) RKHS norm. Restricted to a bounded domain $\Omega\subset\mathbb{R}^{n}$ with Lipschitz boundary, as is of interest in the setting of this paper, the Wendland kernel generates the corresponding Sobolev space $H^{\sigma_{n,l}}(\Omega)$ [24, Theorem 4.1]. Results from kernel approximation theory allow for a more explicit bound on the projection error in terms of the fill distance.
@@
Let $l\in\mathbb{N}$, define $\sigma_{n,l}=\frac{n+1}{2}+l$ and denote by $\mathbb{H}$ the Sobolev space $H^{\sigma_{n,l}}(\Omega)$ induced by the Wendland kernel $\text{\sf{k}}_{n,l}$ as RKHS. Assume that $\mathcal{K}\in L(\mathbb{H},\mathbb{H})$ and $\mathcal{K}\in L(C_{b}(\Omega),C_{b}(\Omega))$. Furthermore, for $\mathcal{X}=\{x_{1},\ldots,x_{d}\}\subset\Omega$, let $m_{j}=m\in\mathbb{N}$ i.i.d. samples from every $\rho_{x_{j}}$, $j\in\{1,\dots,d\}$, be given. Then, for every $\varepsilon>0$ satisfying $m\geq\frac{2C_{2}^{2}}{\varepsilon^{2}}\log(2d)$ and all $f\in\mathbb{H}$ with $\|f\|_{\mathbb{H}}\leq 1$, we have the pointwise error bound
$\left\|(\mathcal{K}-\mathcal{K}^{\mathrm{MC}})f\right\|_{\infty}\leq$ $C_{1}\,(1+\|\mathcal{K}\|_{\mathbb{H}\to\mathbb{H}})\cdot h_{\mathcal{X}}^{l+1/2}+\varepsilon\cdot C_{2}\cdot\sqrt{\max_{v\in\mathbbm{1}}v^{\top}\mathbf{K}_{\mathcal{X}}^{-1}v}$ (22)
with probability at least $1-2d\exp\left(-\frac{m\varepsilon^{2}}{2C_{2}^{2}}\right)$, where $C_{1},C_{2}>0$ are constants only depending on the kernel $\text{\sf{k}}_{n,l}$, and $\mathbbm{1}=\{v\in\mathbb{R}^{d}:|v_{i}|=1\}$.
Note that the bounds of Theorems 2.6 and 2.8 depend on the smallest eigenvalue of the kernel matrix $\mathbf{K}_{\mathcal{X}}$. Tikhonov regularization is a means to control the latter, for the price of introducing a deterministic bias. Section 4 addresses the impact of regularization on the deterministic (Proposition 4.5), probabilistic (Corollary 4.9), and total error bounds (Corollary 4.12).